# A data-driven framework for team selection in Fantasy Premier League

### v2

@@ 8 Experimental results and discussion
Figure 17: Comparing Results with [santorum2025FPL].
We examine the hybrid metric (Eq. (28)) applied to averaging/simulation families (Fig. 9) and to ARIMA and ICT (Fig. 10). Across families, the 1:2 setting (favoring predictions) is more reliable than 2:1 (favoring realized points): it typically provides a small uplift or remains competitive, whereas 2:1 often dilutes the signal by over-weighting historical noise. For ARIMA (0,0,1), hybridization does not improve the season-long trajectory, implying that the ARIMA residual structure already captures the predictable component. For ICT, the hybrid (1:2) variant offers a clear improvement over plain ICT, consistent with the hybrid’s ability to fuse direct non-scoring proxies with recent realized outcomes.
We next vary the starting-XI budget (bench receives the residual to reach £100m) with fixed teams (no weekly repick). Figures 1114 summarize trends across families. For simple averages, tighter starting budgets (down to $\sim£65$-£70m) can outperform the default £83.5m, consistent with stronger benches covering blanks and injuries. For weighted averages, reducing the starting budget generally hurts, suggesting the best performers (recent form) reside among premium or mid-premium assets. In the ARIMA family, ARIMA (0,0,1) is relatively insensitive across £75–£83.5m, while ARIMA (1,0,0) benefits from tighter starting budgets (£70–£80m); tightening the budget leads to a decline in the performance of the ARIMA(1,0,1) model. While the hybrid ICT model does not change much with varying budgets, the ICT score model achieves its highest performance with the default budget; the Monte Carlo model also benefits from a tighter budget, except at £60 million, where talent scarcity dominates.
We now re-pick the starting XI from a fixed 15-man roster each gameweek (no transfers), promoting bench players when their expected points exceed a starter’s and the full roster remains 22-55-55-33 compliant. A rolling policy helps the simple average (notably around £70-£80m) by enabling timely bench promotions; it has little effect on weighted averages and ARIMA (0,0,1). This, however, was not the case in ARIMA (1,0,1), where the rolling setting with £70-£80m outperformed the default model. Figures 12-15 summarize trends across families. For ARIMA (1,0,0), rolling selection is a major uplift, reaching almost $\sim$700 cumulative points with £70–£75m starting budgets, the strongest result in our study. This indicates that a modest trend model benefits disproportionately from active lineup refresh within a fixed squad.
@@
Figure 16 reports the best-performing method within each model family. Table  LABEL:teb:summary provides the final cumulative values for each model in the above-mentioned figure and their comparative rankings. Figure 17 then benchmarks two models from this study, an ARIMA(1,0,0) forecaster with a rolling window (starting-XI budget of £70m) and a weighted-average forecaster, against the four approaches in [santorum2025FPL]. In brief, the Greedy algorithm builds a 15-player squad by ranking a cost-performance (“goodness”) index based on predicted mean season points, with the weekly XI chosen at random; Mean-MIP-random uses Mixed Integer Programming (MIP) on predicted mean scores to select the initial 15, again with a random weekly XI; Mean-MIP-rank keeps the MIP-optimized squad but selects the XI each Gameweek via MIP using adjusted, GW-specific scores (form, home/away, opponent difficulty); and Mean-MIP-online updates the predictions every Gameweek and re-optimizes the XI via MIP while holding the initial squad fixed.
@@
Table 2: Weighted Average Team for Gameweek 27.
@@
Table 3: ARIMA (1,0,0) Rolling (Budget = 70) Team for Gameweek 27
@@
Table 4: Monte Carlo Team for Gameweek 27.
@@
Table 5: Hybrid Simple Average (1:2) Team for Gameweek 27.
@@
Table 6: Linear Regression Team for Gameweek 27.
@@
Table 7: ICT Score Team for Gameweek 27.
@@
As shown in Figure 17, both our ARIMA(1,0,0)-Rolling and weighted-average models outperform the baselines across most of the season. The panel also plots week-by-week differences (ours minus the best competing method); the negative dip in Gameweek 29 stems from data availability rather than model behavior. The only exceptions are Gameweeks 29 and 38, where the MIP-online approach is ahead; otherwise, our two models lead throughout.
Tables 27 present the teams selected by our top model families for Gameweek 27, revealing clear structural regularities across methodologies. Midfield-centric formations, particularly 3-5-2, dominate in almost all approaches, underscoring the models’ collective preference for midfielders’ attacking output and bonus potential under the FPL scoring system. The linear regression model leans toward a 3-4-3 configuration.
Regarding captaincy, Bukayo Saka is the armband choice for five of the six models, highlighting his centrality in the data-driven selection process. Cross-model consistency is also evident in the frequent inclusion of Phil Foden, Cole Palmer, and Ollie Watkins—the latter appearing in every best XI, consistent with Premier League’s Gameweek 27 review (“GW27 stats: Brave FPL managers rewarded for captaining Watkins”). [^7] https://www.premierleague.com/news/3917209 Arsenal assets, especially defenders Gabriel dos Santos Magalhães and William Saliba, recur prominently, reflecting both strong defensive metrics and attacking involvement.
In terms of premium allocation, Mohamed Salah and Erling Haaland are included across several lineups but seldom together, as their combined cost exceeds typical budget constraints unless projected returns justify the expenditure. The goalkeeper position remains the most variable: the ARIMA and Monte Carlo models favor Đorđe Petrović, the weighted average opts for Jordan Pickford, and the hybrid approach selects Alphonse Areola, illustrating contrasting valuation strategies between clean-sheet reliability and save-point potential.
Bench selections are largely consistent, with Conor Bradley and Cameron Archer repeatedly chosen as cost-efficient enablers. Overall, the model outputs exhibit a pronounced convergence toward big teams’ assets(e.g., Arsenal and Liverpool) and Aston Villa, balanced by strategic differentials in defense and goalkeeping designed to optimize both expected value and formation flexibility.
@@ 9 Conclusion
This paper presents an integer–programming framework with a robust variant that turns player-level forecasts into valid FPL lineups. The optimizer selects the starting XI, bench, and captain under budget, formation, and club limits, using data–derived expected points as inputs. To supply these inputs, we compared averaging methods, lightweight time–series forecasting, simulation techniques, and a hybrid approach that blends recent outcomes with feature-based predictions. We also examined the impact of tighter budgets and rolling windows.
Recency-weighted averages and low-order ARIMA models provide strong, stable baselines. Hybrid and robust variants often improve performance, though gains are not universal. The optimizer frequently favors a 3-5-2 shape, reflecting midfielders’ outsized contribution to FPL scoring. Overall, translating noisy forecasts into constraint-aware squads yields transparent, reproducible, and practically useful improvements.
Future work includes embedding transfers and chips in a rolling-horizon setting, allowing week-to-week captaincy, refining goalkeeper objectives, and enriching the feature set with external match-level data. Potential future changes in the FPL rules can also be incorporated into the optimization model.
@@
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
CRediT authorship contribution statement

### v3

@@ 8 Experimental results and discussion
Figure 26: Family/method comparison against the benchmark ARIMA (1,0,0) Rolling (Budget = £70), shown as weekly uplift ($\Delta=$ method-benchmark) over GW 27–38. Each violin summarises the distribution of weekly $\Delta$ values, with points for individual gameweeks. Asterisks on the right margin indicate the significance of a paired Wilcoxon signed-rank test versus zero uplift (BH–FDR adjusted).
Figure 27: Absolute weekly score distributions for the top-10 methods, selected by median weekly uplift relative to the benchmark ARIMA (1,0,0) Rolling (Budget = £70). Each violin shows the distribution of weekly fantasy points over GW 27–38, with dots marking individual weeks.
Figure 28: Comparing results with (santoro2025FPL).
Figure 25 reports the best-performing method within each model family over Gameweeks 27–38. The curves show that the ARIMA(1,0,0) rolling forecaster with a £70m starting–XI budget achieves the highest cumulative score, followed closely by the weighted-average model, while the Monte Carlo and hybrid simple-average approaches trail behind. Table LABEL:tab:summary complements this figure by listing the final cumulative scores for all models shown and their comparative rankings.
To further analyze the stability of these results, Figure 26 compares each modeling family to the benchmark ARIMA (1,0,0) Rolling (£70) in terms of weekly points uplift over GW 27–38. This specific ARIMA configuration is adopted as the baseline because, as established in our cumulative performance analysis, it yields the most consistent out-of-sample results. The vertical dashed line at 0 marks parity with the benchmark; violins lying to the right correspond to methods that typically outscore it, while those centered near 0 behave similarly. ICT-based optimization and the linear-regression family deliver clear positive gains, identifying them as strong alternatives for weekly upside. In contrast, Exponential smoothing and its hybrid variant (Hybrid ExpSm 1:2) show the largest negative uplifts, with almost all weekly differences falling below zero. Monte Carlo, non-rolling ARIMA variants, and the weighted-average baselines cluster tightly around zero (or slightly negative), indicating performance comparable to, but generally outpaced by the benchmark once multiple testing is controlled for.
Figure 27 zooms in on the ten best strategies identified in the uplift analysis (Figure 26) and presents their absolute weekly scores rather than differences from the benchmark. Each horizontal violin summarizes the distribution of points across GW 27–38, with dots marking individual gameweeks. The top-ranked Hybrid ARIMA 1:2 exhibits the right-most distribution, characterized by a high median and few low-scoring weeks, indicating consistently strong performance. The various rolling ARIMA (0,0,1) and ARIMA (1,0,0) configurations cluster closely together with very similar medians and spreads, suggesting that minor budget tweaks mainly shift scores slightly without altering variability. Simple Avg Rolling also attains competitive medians but exhibits a somewhat wider violin, reflecting greater week-to-week volatility. Overall, the figure shows that the leading approaches differ by modest right-shifts in their entire score distributions rather than by occasional extreme gameweeks.
For external benchmarking, we concentrate on two strong yet relatively simple models—ARIMA(1,0,0) with a rolling window and a recency-weighted average forecaster—because they offer a good trade-off between performance and interpretability, while avoiding the extra tuning overhead of hybrid or ICT-based objectives.
@@
Figure 28 then benchmarks these two models against the four approaches proposed in santoro2025FPL. The Greedy algorithm constructs a 15-player squad by ranking a cost–performance (“goodness”) index based on predicted mean season points, with the weekly XI chosen at random. Mean-MIP-random uses Mixed Integer Programming (MIP) on predicted mean scores to select the initial 15, again with a random weekly XI; Mean-MIP-rank keeps the MIP-optimized squad but chooses the XI each Gameweek via MIP using adjusted, GW-specific scores (form, home/away status, opponent difficulty); and Mean-MIP-online updates predictions every Gameweek and re-optimizes the XI via MIP while holding the initial squad fixed. As shown in the upper panel, our ARIMA(1,0,0)-rolling and weighted-average models outperform these baselines for most of the season. The lower panel plots the week-by-week score difference between our best model and the strongest competing method; the negative dip in Gameweek 29 is driven by data availability rather than model behavior. The only Gameweeks in which the Mean-MIP-online approach is ahead are 29 and 38; otherwise, our two models maintain the lead throughout.
@@
(a) Weighted Average Team for Gameweek 27.
@@
(b) ARIMA (1,0,0) Rolling (Budget = 70) Team for Gameweek 27.
@@
(a) Monte Carlo Team for Gameweek 27.
@@
(b) Hybrid Simple Average (1:2) Team for Gameweek 27.
@@
(a) Linear Regression Team for Gameweek 27.
@@
(b) ICT Score Team for Gameweek 27.
@@
Tables 2(a)4(b) list the Gameweek 27 teams produced by our forecasting models (weighted average, ARIMA, Monte Carlo, and hybrid simple average) together with two additional baselines (linear regression and ICT score), revealing structural regularities in the resulting squads. Midfield-centric formations, particularly 3–5–2, appear in five of the six approaches, indicating a shared preference for midfielders’ attacking output and bonus potential under the FPL scoring system. The linear regression model is the only exception, leaning toward a 3–4–3 configuration.
Regarding captaincy, Bukayo Saka is selected as captain in four of the six models, underlining his central role in the data-driven selection process during that season. Cross-model consistency is also evident in the frequent inclusion of Phil Foden, Cole Palmer, and Ollie Watkins—the latter appearing in every best XI, in line with the Premier League’s Gameweek 27 review (“GW27 stats: Brave FPL managers rewarded for captaining Watkins”). [^7] https://www.premierleague.com/news/3917209 Arsenal assets, especially defenders Gabriel dos Santos Magalhães and William Saliba, recur prominently, reflecting a combination of strong defensive metrics and attacking involvement.
In terms of premium allocation, Mohamed Salah and Erling Haaland are each included in several line-ups but do not appear together in any of the selected squads, as their combined cost would require substantial sacrifices elsewhere unless projected returns clearly justify the investment. The goalkeeper position exhibits the greatest variability: the ARIMA and Monte Carlo models favour Đorđe Petrović, the weighted-average model opts for Jordan Pickford, the hybrid model selects Alphonse Areola, while the linear-regression and ICT-score teams choose Ivo Grbić and James Trafford, respectively. Bench selections are nonetheless broadly consistent, with Conor Bradley and Cameron Archer repeatedly appearing as low-cost enablers. Overall, the outputs converge on assets from leading clubs (notably Arsenal and Liverpool) and Aston Villa, complemented by strategic differentials in defence and goalkeeping to balance expected value with formation flexibility.
@@ 9 Conclusion
This paper introduced an integer-programming framework, together with a robust variant, for turning player-level forecasts into valid FPL lineups. The optimizer selects the starting XI, bench, and captain under budget, formation, and club constraints, using data-derived expected points as inputs. To generate these inputs, we compared averaging methods, lightweight time-series forecasting, simulation techniques, and a hybrid approach that blends recent outcomes with feature-based predictions, and we examined the impact of tighter budgets and rolling windows.
Across our experiments, recency-weighted averages and low-order ARIMA models provided strong and stable baselines. Hybrid variants often improved performance, while robust variants helped for ICT-based objectives but did not consistently outperform their non-robust counterparts. When benchmarked against the four strategies of Santorosantoro2025FPL, our ARIMA(1,0,0) rolling and weighted-average models delivered higher cumulative scores over Gameweeks 27–38, indicating that relatively simple forecasts can outperform more elaborate pipelines once embedded in an appropriate integer-programming formulation. The optimizer frequently favored a 3–5–2 shape, reflecting midfielders’ relatively high marginal contribution under the FPL scoring rules. Overall, translating noisy forecasts into constraint-aware squads produced transparent, reproducible, and practically useful improvements over benchmark strategies.
Future work includes embedding transfers and chips in a rolling-horizon setting, allowing week-to-week captaincy decisions, refining goalkeeper objectives, and enriching the feature set with external match-level data. Potential changes in FPL rules can also be incorporated directly into the optimization model.
@@
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Author contribution
@@
Data Availability Statement
The datasets and source code are available at the link provided in the paper.
Funding Information
Not Applicable
Research Involving Human and /or Animals
Not Applicable
Informed Consent
Not Applicable